# Exploring the Relationship between Analogy Identification and Sentence Structure Encoding in Large Language Models

### v1

@@ Abstract
Identifying analogies plays a pivotal role in human cognition and language proficiency. In the last decade, there has been extensive research on word analogies in the form of “A is to B as C is to D.” However, there is a growing interest in analogies that involve longer text, such as sentences and collections of sentences, which convey analogous meanings. While the current NLP research community evaluates the ability of Large Language Models (LLMs) to identify such analogies, the underlying reasons behind these abilities warrant deeper investigation. Furthermore, the capability of LLMs to encode both syntactic and semantic structures of language within their embeddings has garnered significant attention with the surge in their utilization. In this work, we examine the relationship between the abilities of multiple LLMs to identify sentence analogies, and their capacity to encode syntactic and semantic structures. Through our analysis, we find that analogy identification ability of LLMs is positively correlated with their ability to encode syntactic and semantic structures of sentences. Specifically, we find that the LLMs which capture syntactic structures better, also have higher abilities in identifying sentence analogies.
@@ 1 Introduction
Analogies facilitate the transfer of meaning and knowledge from one domain to another. Making and identifying analogies is a central tenet in human cognition Hofstadter (2001); Holyoak et al. (2001) and is aided by humans’ ability to process the structure of language. In the domain of NLP, several types of textual analogies are identified, such as word analogies Yuan et al. (2023); Gladkova et al. (2016); Gao et al. (2014), proportional word analogies Chen et al. (2022); Ushio et al. (2021); Szymanski (2017); Drozd et al. (2016), sentence-analogies Afantenos et al. (2021); Zhu and de Melo (2020); Wang and Lepage (2020) and more recently analogies between procedural/long text Sultan and Shahaf (2022).
Figure 1: This pipeline details the process of quantifying the LLMs abilities to capture sentence structure via SyntScore and SemScore scores for a given sentence. In this work, we apply this process to a dataset of 100K sentences. The dataset is divided into 0.8 for training the structure probe and 0.1 for testing.
Despite the existence of several established benchmarks (e.g., SuperGLue and GLUE) which evaluate the abilities of LLMs extrinsically, Wijesiriwardene et al. (2023) propose a more challenging intrinsic benchmark that focuses on LLMs’ ability to identify analogies across a range of complexities. Researchers argue that language understanding necessitates humans to grasp the inherent structure of language, including parse trees and logical formulas Chomsky (2002); Schank (1972). Furthermore, the identification of analogies relies on the utilization of implicit relational knowledge embedded within the relational structure of language Gentner (1983). Consequently, the present study investigates the potential relationship between LLMs’ ability to identify analogies and their aptitude for encoding sentence structures.
This work looks explicitly at sentence-level analogies  [^1] An analogy in the context of our manuscript always means sentence-analogy., which is a clear starting point for this work. We also look at LLMs’ abilities to capture both syntactic and semantic structures of sentences. Specifically, our main contribution is an empirical analysis of the correlation between the analogy identification ability and sentence structure encoding abilities of LLMs. Additionally, we extend the sentence structure probing techniques introduced by Hewitt and Manning (2019) (which only supports BERT and ELMo) to further work with encoder-decoder-based LLMs and LLMs that use two transformer architectures. Finally, we utilize the structure probing technique originally used for syntactic structure probing in the novel context of semantic structure probing.
@@ 2 Related Work
Compared to syntactic parsing, the NLP communities’ interest in semantic parsing is growing. Semantic parsing maps natural language sentences to a complete, formal meaning representation. Semantic parsing is achieved via combining the Semantic Role Labelling (SRL) approaches with syntactic dependency parsing Hajic et al. (2009); Surdeanu et al. (2008) and more recently via semantic dependency parsing Oepen et al. (2014, 2015). This work uses the semantic dependency parsing approach based on mean field variational inference (MFVI) augmented with character and lemma level embeddings introduced by Wang et al. (2019).
@@ 3.1 Dataset
We experiment on a dataset of 100K English sentences. Specifically, the dataset used in this work is randomly picked from the sentence corpus of levels three, four and five of the analogy taxonomy introduced in Wijesiriwardene et al. (2023). The composition of the dataset is presented in Table 1 (duplicates removed).
@@ 3.2 Large Language Models and their Ability to Capture Sentence Analogies
We experiment on the eight language models used in a study by Wijesiriwardene et al. (2023) namely, BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), ALBERT Lan et al. (2019), LinkBERT Yasunaga et al. (2022), SpanBERT Joshi et al. (2020) and XLNet Yang et al. (2019) which are encoder-based LLMs, T5 Raffel et al. (2020), an encoder-decoder-based LLM and ELECTRA Clark et al. (2020), an LLM based on two transformer architectures. We refer readers to cited publications for details on the LLMs used in this work.
@@ 3.3 Large Language Models and their Ability to Capture Sentence Structures
The authors introduce a path distance metric and a path depth metric for evaluation. The distance metric captures the path length between each pair of words measured by Undirected Unlabeled Attachment Score (UUAS) and average Spearman correlation of true to predicted distances (DSpr). The depth metric evaluates the model’s ability to identify the sentence’s root, measured as root accuracy percentage. Additionally, the depth metric also evaluates the ability of the model to recreate the word order based on their depth in the parse tree identified as Norm Spearman (NSpr.) [^2] We do not use NSpr. in this work. We refer the readers to Hewitt and Manning (2019) for further details on the technique and evaluation metrics.
@@ 4 Experimental Setup
To obtain the SemScore (see Figure 1), we first parse all the sentences in our dataset using the MFVI approach Wang et al. (2019). The resulting semantically parsed sentences (in CoNLL-U format) [^3] https://universaldependencies.org/format.html and the LLM embeddings of the original sentences are sent to the structure probing technique Hewitt and Manning (2019). The structure probe is trained on 80K sentences from the dataset and results, DSpr and UUAS representing parse distance and root accuracy (RootAcc) representing parse depth are reported on the test split with 10K sentences. Finally, the SemScore is computed as a combined score by taking the mean of the z-score normalizations of these three measures $Z_{DSpr},Z_{UUAS},Z_{RootAcc}$ (see Table 2).
@@
To obtain the SyntScore (see Figure 1), we follow the same steps but parse the sentences syntactically.
Finally, we calculate the Spearman’s rank correlation (SRC) and Kendall’s rank correlation (KRC) between AnalogyScore and SyntScore, as well as AnalogyScore and SemScore.
@@ 4.1 Implementation Details
When extending the structure probing technique by Hewitt and Manning (2019) to facilitate additional LLMs, we use the HuggingFace implementation [^4] https://huggingface.co/models of the LLMs. For semantic parsing, we use the trained mean field variational inference (MFVI) model augmented with character and lemma-level embeddings provided by the SuPar [^5] https://github.com/yzhangcs/parser package. For syntactic parsing of the sentences we employ Stanford CoNLL-U dependency parser [^6] https://nlp.stanford.edu/software/nndep.html.
@@ 5.2 Analogy Identification and Syntactic Structure Encoding Abilities of LLMs
We use SRC and KRC values to analyze the correlation between LLMs’ ability to identify sentence analogies denoted by AnalogyScore and LLMs’ ability to encode syntactic structures of sentences denoted by SyntScore. Both correlation measures show a positive correlation between AnalogyScore and SyntScore. Specifically, the SRC between AnalogyScore and SyntScore is 0.95 ($p<0.001$). The KRC between AnalogyScore and SyntScore is 0.86 ($p=0.002$).
@@ 6 Limitations
Even though Abstract Meaning Representation (AMR) Banarescu et al. (2013) is a popular and widely used technique to parse sentence meanings, in this work, we use MFVI, a semantic parsing approach introduced by Wang et al. (2019) because of the limitations posed by the structure probing technique used Hewitt and Manning (2019). This technique requires the mapped LLM embeddings and semantic dependency parsed sentences to be the same size. However, as it is known, AMRs abstract away from the syntactic idiosyncrasies of the language and overlook certain auxiliary words from the parse results, limiting its use in this work.
The present study employs a semantic parsing technique reported to exhibit a high accuracy level of 94% (Wang, 2019). However, it is important to note that for the purposes of our investigation, we make the assumption that the semantically parsed sentences generated by this particular method are entirely accurate, thereby employing them as the gold standard data. It is worth mentioning that this choice may introduce some degree of bias into our examination of the semantic structure probing.
We obtain sentence-analogy pairs provided by Wijesiriwardene et al. (2023) [^7] We thank the authors of Wijesiriwardene et al. (2023) for providing access to the datasets. and split the pairs to obtain single sentences. However, it should be noted that the subset of sentences employed in this work does not exhibit a balanced distribution across the three levels of the taxonomy.
7 Conclusion and Future Avenues
This work explores the correlation between LLMs’ ability to identify analogies and encode sentence structures in their embeddings. Through detailed experiments, we show that the analogy identification ability of LLMs is positively correlated with their ability to encode syntactic and semantic structures. Particularly, LLMs that better capture syntactic structures have a higher correlation to analogy identification. As a next step, we would like to assess the possibility of extending this approach in the context of explainability of LLMs.

### v3

# On the Relationship between Sentence Analogy Identification and Sentence Structure Encoding in Large Language Models

@@ Abstract
The ability of Large Language Models (LLMs) to encode syntactic and semantic structures of language is well examined in NLP. Additionally, analogy identification, in the form of word analogies are extensively studied in the last decade of language modeling literature. In this work we specifically look at how LLMs’ abilities to capture sentence analogies (sentences that convey analogous meaning to each other) vary with LLMs’ abilities to encode syntactic and semantic structures of sentences. Through our analysis, we find that LLMs’ ability to identify sentence analogies is positively correlated with their ability to encode syntactic and semantic structures of sentences. Specifically, we find that the LLMs which capture syntactic structures better, also have higher abilities in identifying sentence analogies.
On the Relationship between Sentence Analogy Identification and Sentence Structure Encoding in Large Language Models
Thilini Wijesiriwardene1thanks:   Corresponding author , Ruwan Wickramarachchi1, Aishwarya Naresh Reganti2 Vinija Jain3,4thanks:   Work does not relate to position at Amazon.  , Aman Chadha3,4footnotemark:   , Amit Sheth1, Amitava Das1 1AI Institute, University of South Carolina, USA, 2Carnegie Mellon University, Pittsburgh, USA, 3Amazon GenAI, USA, 4Stanford University, USA thilini@sc.edu
@@ 1 Introduction
Analogies facilitate the transfer of meaning and knowledge from one domain to another. Making and identifying analogies is a central tenet in human cognition Hofstadter (2001); Holyoak et al. (2001) and is aided by humans’ ability to process the structure of language. In the domain of NLP, several types of textual analogies are identified, such as word analogies Yuan et al. (2023); Gladkova et al. (2016); Gao et al. (2014), proportional word analogies Chen et al. (2022); Ushio et al. (2021); Szymanski (2017); Drozd et al. (2016), sentence-analogies Afantenos et al. (2021); Zhu and de Melo (2020); Wang and Lepage (2020) and more recently analogies of procedural/long text Sultan and Shahaf (2022). This work explicitly looks at sentence-level analogies which are sentence pairs that are analogues in meaning to each other  [^1] For more details on sentence analogies please refer to Wijesiriwardene et al. (2023).
Figure 1: This pipeline details the process of quantifying the LLMs abilities to capture sentence structure via SyntScore and SemScore values for a given sentence. In this work, we apply this process to a dataset of 100K sentences. The dataset is divided into 0.8 for training the structure probe and 0.1 for testing.
Despite the existence of several established benchmarks (e.g., SuperGLUE Wang et al. (2019a) and GLUE Wang et al. (2018)) which evaluate the abilities of LLMs extrinsically, Wijesiriwardene et al. (2023) propose a more challenging intrinsic benchmark that focuses on LLMs’ ability to identify analogies across a range of complexities.
Identification of analogies relies on the utilization of implicit relational knowledge embedded within the relational structure of language Gentner (1983).
In this work we aim to explore the relationship between sentence analogy identification abilities and syntactic and semantic structure encoding abilities of LLMs [^2] Our code is available at: https://github.com/Thiliniiw/llms-synt-struct-sentence-analogies.
Specifically, our main contribution is an analysis of the relationship between the analogy identification ability and sentence structure encoding abilities of LLMs. Additionally, we extend the sentence structure probing techniques introduced by Hewitt and Manning (2019) (which only supports BERT and ELMo) to further work with encoder-decoder-based LLMs and LLMs that use two transformer architectures. Finally, we extend the structure probing technique originally used for syntactic structure probing in the novel context of semantic structure probing.
@@ 2 Related Work
Compared to syntactic parsing, the NLP communities’ interest in semantic parsing is growing. Semantic parsing maps natural language sentences to a complete, formal meaning representation. Semantic parsing is achieved via combining the Semantic Role Labelling (SRL) approaches with syntactic dependency parsing Hajic et al. (2009); Surdeanu et al. (2008) and more recently via semantic dependency parsing Oepen et al. (2014, 2015). This work uses the semantic dependency parsing approach based on mean field variational inference (MFVI) augmented with character and lemma level embeddings introduced by Wang et al. (2019b).
@@ 3.1 Dataset
We experiment on a dataset of 100K English sentences. Specifically, the dataset used in this work is randomly picked from the sentence corpus of levels three, four and five of the analogy taxonomy introduced in Wijesiriwardene et al. (2023). The composition of the dataset is presented in Table 1 (duplicates removed). Specifically, we obtain sentence-analogy pairs provided by Wijesiriwardene et al. (2023) and split the pairs to obtain single sentences used in this work.
@@ 3.2 Large Language Models and their Ability to Capture Sentence Analogies
We experiment on the eight language models used in a study by Wijesiriwardene et al. (2023) namely, BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), ALBERT Lan et al. (2019), LinkBERT Yasunaga et al. (2022), SpanBERT Joshi et al. (2020) and XLNet Yang et al. (2019) which are encoder-based LLMs, T5 Raffel et al. (2020), an encoder-decoder-based LLM and ELECTRA Clark et al. (2020), an LLM based on two transformer architectures. We refer readers to cited publications for details on the specific LLMs.
@@ 3.3 Large Language Models and their Ability to Capture Sentence Structures
The authors introduce a path distance metric and a path depth metric for evaluation. The distance metric captures the path length between each pair of words measured by Undirected Unlabeled Attachment Score (UUAS) and average Spearman correlation of true to predicted distances (DSpr). The depth metric evaluates the model’s ability to identify a sentence’s root, measured as root accuracy percentage. Additionally, the depth metric also evaluates the ability of the model to recreate the word order based on their depth in the parse tree identified as Norm Spearman (NSpr.) [^3] We do not use NSpr. in this work. We refer the readers to Hewitt and Manning (2019) for further details on the technique and evaluation metrics.
@@ 4 Experimental Setup
To obtain the SemScore (see Figure 1), we first parse all the sentences in our dataset using the MFVI approach Wang et al. (2019b). The resulting semantically parsed sentences (in CoNLL-U format) [^4] https://universaldependencies.org/format.html and the LLM embeddings of the original sentences are then sent to the structure probing technique Hewitt and Manning (2019). The structure probe is trained on 80K sentences from the dataset and the DSpr and UUAS values representing parse distance and root accuracy (RootAcc) value representing parse depth are reported on the test split with 10K sentences. Finally, the SemScore is computed as a combined score by taking the mean of the z-score normalizations of these three measures $Z_{DSpr},Z_{UUAS},Z_{RootAcc}$ (see Table 2).
@@
To obtain the SyntScore (see Figure 1), we follow the same steps but parse the sentences syntactically. Finally, we calculate the Spearman’s rank correlation (SRC) and Kendall’s rank correlation (KRC) between AnalogyScore and SyntScore, as well as AnalogyScore and SemScore.
@@ 4.1 Implementation Details
When extending the structure probing technique by Hewitt and Manning (2019) to facilitate additional LLMs, we use the HuggingFace implementation [^5] https://huggingface.co/models of the LLMs. For semantic parsing, we use the trained mean field variational inference (MFVI) model augmented with character and lemma-level embeddings provided by the SuPar [^6] https://github.com/yzhangcs/parser. For syntactic parsing of the sentences we employ Stanford CoNLL-U dependency parser [^7] https://nlp.stanford.edu/software/nndep.html.
@@
Model AnalogyScore SyntScore SemScore
Score Rank Score Rank Score Rank
AlBERT 0.645 7 -2.14 8 -0.19 5
BERT 0.505 3 0.68 3 0.14 3
Electra 0.516 4 0.66 4 -0.21 6
LinkBERT 0.608 6 0.22 5 -0.56 8
RoBERTa 0.458 1 0.78 1 0.84 1
SpanBERT 0.461 2 0.72 2 -0.02 4
T5 0.524 5 -0.27 6 -0.32 7
XLNet 0.747 8 -0.64 7 0.32 2
@@ 5 Results
In this section, we look at the findings of this work with regard to semantic and syntactic structure encoding abilities and analogy identification abilities of LLMs.
@@ 5.2 Analogy Identification and Syntactic Structure Encoding Abilities of LLMs
We use SRC and KRC values to analyze the correlation between LLMs’ ability to identify sentence analogies denoted by AnalogyScore and LLMs’ ability to encode syntactic structures of sentences denoted by SyntScore. Both correlation measures show a significant positive correlation between AnalogyScore and SyntScore. Specifically, the SRC between AnalogyScore and SyntScore is 0.95 ($p<0.001$). The KRC between AnalogyScore and SyntScore is 0.86 ($p=0.002$).
@@ 6 Limitations
Several contemporary probing techniques, such as those outlined in Voita and Titov (2020) and Pimentel et al. (2020), have emerged subsequent to the methodology employed in the present investigation Hewitt and Manning (2019). Nevertheless, in the context of our current study, we have only chosen to employ Hewitt and Manning (2019) owing to its adaptable nature, which facilitates extension to various LLMs that are of particular interest to our current research.
Even though Abstract Meaning Representation (AMR) Banarescu et al. (2013) is a popular and widely used technique to parse sentences semantically, in current work, we use MFVI, a semantic parsing approach introduced by Wang et al. (2019b) because of the limitations posed by the structure probing technique used Hewitt and Manning (2019). This technique requires the mapped LLM embeddings and semantic dependency parsed sentences to be of the same length. However, as it is known, AMRs abstract away from the syntactic idiosyncrasies of the language and overlook certain auxiliary words from the parse results, limiting its use in this work.
The present study employs a semantic parsing technique reported to exhibit a high accuracy level of 94% Wang et al. (2019b). However, it is important to note that for the purposes of our investigation, we make the assumption that the semantically parsed sentences generated by this particular method are entirely accurate, thereby employing them as the gold standard data. It is worth mentioning that this choice may introduce some degree of bias into our examination of the semantic structure probing.
7 Conclusion and Future Directions
This work explores the relationship between LLMs’ ability to identify sentence analogies and encode sentence structures in their embeddings. Through detailed experiments, we show that the sentence analogy identification ability of LLMs is positively correlated with their ability to encode syntactic and semantic structures of sentences. Particularly, LLMs that better capture syntactic structures have a higher correlation to analogy identification. In summary this work explores how LLMS utilize the knowledge of semantic and syntactic structures of sentences to identify analogies. Moving forward, we aim to explore the potential of extending the current approach to enhance explainability of LLMs within the broader domain of NLP.
Acknowledgements
We thank the anonymous reviewers for their constructive comments. This work was supported in part by the NSF grant #2335967: EAGER: Knowledge-guided neurosymbolic AI with guardrails for safe virtual health assistants. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding organization.